OpenStudy (i_love_my_nieces):

Help needed! What is the rule for the function shown in the table? 1.) y = 1/3x + 1 2.) y = 3x + 1 3.) y = 1/3x 4.) y = 1/2x + 1

OpenStudy (i_love_my_nieces):

@Michele_Laino

OpenStudy (lolipop_drum312):

you get it?

OpenStudy (i_love_my_nieces):

Yes

OpenStudy (lolipop_drum312):

ok good

OpenStudy (i_love_my_nieces):

Is s/he right?

OpenStudy (michele_laino):

we have to start from this formula: y=a*x+b and we have to determine the coefficients a, and b, using data stored into your table

OpenStudy (michele_laino):

now I consider the first point, namely: x=0, and y=1

OpenStudy (michele_laino):

substituting those value, I get: 1=a*0+b what is b?

OpenStudy (michele_laino):

values*

OpenStudy (michele_laino):

@I_love_my_nieces what is b, please?

OpenStudy (i_love_my_nieces):

How would I figure out what b is

OpenStudy (michele_laino):

here is my relationship, after that substitution: 1= a*0+ b

OpenStudy (lolipop_drum312):

i think B is correct

OpenStudy (i_love_my_nieces):

But you just your are 100% sure that c is the answer...

OpenStudy (lolipop_drum312):

sorry i typed the wrong i met to say B

OpenStudy (lolipop_drum312):

the answer letter is just differ but its B

OpenStudy (i_love_my_nieces):

So my answer is y=3x+1?

OpenStudy (i_love_my_nieces):

@Lolipop_Drum312 @Michele_Laino

OpenStudy (michele_laino):

that's right! @I_love_my_nieces

OpenStudy (lolipop_drum312):

correct!!